# Emulation of Complex Matrix Multiplication based on the Chinese Remainder Theorem

### v1

@@ IV-B Performance
Figure 6: Throughput performance in TFLOPS of single-precision complex matrix multiplication on NVIDIA GH200 Grace Hopper Superchip, CUDA Toolkit 13.0.88, and gcc 11.5.0, for $m=k=n$. Methods shown in red italic achieve CGEMM-level accuracy.
@@
Figure 7: Throughput performance in TFLOPS of single-precision complex matrix multiplication on NVIDIA B200 Blackwell GPU with Intel Xeon Platinum 8570 CPU, CUDA Toolkit 12.8.93, and g++ 13.3.0, for $m=k=n$. Methods shown in red italic achieve CGEMM-level accuracy.
@@
Figure 8: Throughput performance in TFLOPS of single-precision complex matrix multiplication on NVIDIA GeForce RTX 5080 GPU with AMD Ryzen 7 3700X CPU, CUDA Toolkit 13.0.88, and gcc 11.5.0, for $m=k=n$. Methods shown in red italic achieve CGEMM-level accuracy.
@@
Figure 9: Throughput performance in TFLOPS of single-precision complex matrix multiplication on AMD Instinct MI300X GPU with AMD EPYC 9534 CPU, ROCm 6.4.2, and clang 19.0.0git, for $m=k=n$. Methods shown in red italic achieve CGEMM-level accuracy.
@@
Figure 10: Throughput performance in TFLOPS of double-precision complex matrix multiplication on NVIDIA GH200 Grace Hopper Superchip, CUDA Toolkit 13.0.88, and gcc 11.5.0, for $m=k=n$. Methods shown in red italic achieve ZGEMM-level accuracy.
@@
Figure 11: Throughput performance in TFLOPS of double-precision complex matrix multiplication on NVIDIA B200 Blackwell GPU with Intel Xeon Platinum 8570 CPU, CUDA Toolkit 12.8.93, and g++ 13.3.0, for $m=k=n$. Methods shown in red italic achieve ZGEMM-level accuracy.
@@
Figure 12: Throughput performance in TFLOPS of double-precision complex matrix multiplication on NVIDIA GeForce RTX 5080 GPU with AMD Ryzen 7 3700X CPU, CUDA Toolkit 13.0.88, and gcc 11.5.0, for $m=k=n$. Methods shown in red italic achieve ZGEMM-level accuracy.
@@
Figure 13: Throughput performance in TFLOPS of double-precision complex matrix multiplication on AMD Instinct MI300X GPU with AMD EPYC 9534 CPU, ROCm 6.4.2, and clang 19.0.0git, for $m=k=n$. Methods shown in red italic achieve ZGEMM-level accuracy.
@@
These performance characteristics correlate strongly with the architectural balance of arithmetic throughput on each GPU. The GH200 provides a ratio of $\text{FP64}:\text{FP32}:\text{INT8}=1:1:29.5$, indicating that INT8 operations are moderately more abundant than FP64 and FP32 units. The B200 exhibits an even more pronounced imbalance, $1:2:121.6$, where INT8 throughput overwhelmingly dominates the floating-point pipelines. This architectural skew directly amplifies the benefit of the proposed kernels, which replace large portions of floating-point arithmetic with INT8-intensive computation. In contrast, the MI300X offers a more subdued ratio of $1:1:16$, providing substantially less INT8 advantage relative to its floating-point throughput. As a result, the computational substitution enabled by our method yields only marginal acceleration on the MI300X, consistent with the limited speedups observed in the benchmarks. For the RTX 5080, the observed performance gains, especially the very large speedups in ZGEMM, suggest a much stronger favoring of INT8 throughput relative to floating-point units. The architectural balance on RTX 5080 corresponds to $\text{FP64}:\text{FP32}:\text{INT8}=1:64:512$, which explains the pronounced acceleration when using INT8-based emulation.
@@
Compared with the Ozaki-I emulation implemented in cuBLAS, the proposed method delivered substantially higher throughput on both GH200 and RTX 5080 GPUs. On the GH200, Ozaki-I with 7–9 slices achieved 25–39 TFLOPS, whereas the proposed method attained 99–136 TFLOPS. On the RTX 5080, Ozaki-I achieved 6.4–10 TFLOPS, while the proposed method reached 19–26 TFLOPS. These disparities arise from the fundamental algorithmic differences: Ozaki-I with $S$ slices requires $S(S+1)/2$ INT8 complex matrix multiplications, whereas the proposed technique with $N$ moduli requires only $N$ INT8 complex matrix multiplications and and further reduces their cost by a factor of $3/4$ using the Karatsuba method. This reduction in both the number and cost of INT8 matrix multiplication kernel invocations directly explains the substantially higher performance of the proposed method.
IV-C Supplemental Comparison for Real-Valued Emulation
@@
Here, we focus on double-precision real matrix multiplication. In this section, we focus on double-precision real matrix multiplication. We compare two emulation strategies: the Ozaki-I based emulation provided by cuBLAS and the Ozaki-II based emulation proposed in [uchino_ozaki2], enhanced with the blocking technique introduced in Section III. This supplemental analysis aims to isolate the behavior of Ozaki-type schemes in the real domain and to clarify the performance implications of incorporating blocking into Ozaki-II under conditions consistent with real-valued DGEMM emulation.
@@
According to [uchino_ozaki2], the fast mode and the accurate mode require 14–18 and 14–17 moduli, respectively, to achieve DGEMM-level accuracy. On the GH200 with matrix size $m=n=k=16384$, the Ozaki-II fast-$N$ configuration with $14\leq N\leq 18$ achieved 63–82 TFLOPS, while the accu-$N$ configuration with $14\leq N\leq 17$ reached 63–75 TFLOPS. Incorporating the blocking technique improved performance across all settings: the blocked fast-$N$ version attained 72–93 TFLOPS, and the blocked accu-$N$ version delivered 71–85 TFLOPS. cuBLAS’s Ozaki-I implementation with 7–10 slices achieved only 20–39 TFLOPS, and the standard DGEMM kernel sustained 61 TFLOPS.
@@
On the RTX 5080 with $m=n=k=8192$, the fast-$N$ and accu-$N$ configurations of the proposed method reached 12–16 TFLOPS and 12–15 TFLOPS, respectively, while cuBLAS’s Ozaki-I achieved 5.4–10 TFLOPS. In comparison, the native DGEMM kernel delivered 0.84 TFLOPS, highlighting the substantial advantage of INT8-intensive Ozaki-type emulation on this consumer architecture.
@@
These results demonstrate that blocking substantially enhances the performance of Ozaki-II for large-scale problems, particularly on bandwidth-rich GPUs such as the H200, and Ozaki-II, whether blocked or unblocked, consistently outperforms cuBLAS’s Ozaki-I due to its reduced number of INT8 matrix multiplication invocations.
@@ V Conclusion
This study proposed emulation methods of complex matrix multiplication based on the Ozaki-II scheme and developed a portable library that supports execution on both AMD and NVIDIA GPUs. The numerical results demonstrated that the proposed methods outperform the standard GEMM routines in cuBLAS and hipBLAS, which are implemented using native FP32 and FP64 arithmetic. In addition, the proposed methods outperform the emulation using the Ozaki-I scheme implemented in cuBLAS.
Future work includes comparisons with the BF16x9 CGEMM emulation algorithm provided for NVIDIA B200 and B300 GPUs in cuBLAS 12.9 and later. A major limitation of the proposed method, as well as emulation-based approaches in general, is the substantial working memory required. This overhead is currently unavoidable when aiming to achieve high performance without relying on FP32 or FP64 matrix multiplication. Addressing this issue remains an open challenge not only for emulation techniques but also for HPC applications that employ them.
The Ozaki-I and Ozaki-II schemes utilize INT8 matrix engines, benefiting from the fact that the error-free low-precision matrix multiplications appearing in these schemes do not require exponent handling, enabling efficient synergy with INT8 operations with INT32 accumulation. When using FP16 or BF16 matrix engines with FP32 accumulation, the effective representable integer range is limited to 24 bits due to the FP32 significand, leading to significantly reduced computational efficiency compared with INT32 accumulation. However, some recent GPUs, such as the NVIDIA B300, provide reduced INT8 capabilities. Supporting such hardware architectures represents an additional direction for future work.

### v3

@@ IV-B Performance
Figure 6: Throughput performance (in TFLOPS) of single-precision complex matrix multiplication on NVIDIA GH200 Grace Hopper Superchip (CUDA Toolkit 13.0.88, gcc 11.5.0) for $m=k=n$. Methods shown in red and italics achieve CGEMM-level accuracy.
@@
Figure 7: Throughput performance (in TFLOPS) of single-precision complex matrix multiplication on NVIDIA B200 Blackwell GPU with Intel Xeon Platinum 8570 CPU (CUDA Toolkit 12.8.93, g++ 13.3.0) for $m=k=n$. Methods shown in red and italics achieve CGEMM-level accuracy.
@@
Figure 8: Throughput performance (in TFLOPS) of single-precision complex matrix multiplication on NVIDIA GeForce RTX 5080 GPU with AMD Ryzen 7 3700X CPU (CUDA Toolkit 13.0.88, gcc 11.5.0) for $m=k=n$. Methods shown in red and italics achieve CGEMM-level accuracy.
@@
Figure 9: Throughput performance (in TFLOPS) of single-precision complex matrix multiplication on AMD Instinct MI300X GPU with AMD EPYC 9534 CPU (ROCm 6.4.2, clang 19.0.0git) for $m=k=n$. Methods shown in red and italics achieve CGEMM-level accuracy.
@@
Figure 10: Throughput performance (in TFLOPS) of double-precision complex matrix multiplication on NVIDIA GH200 Grace Hopper Superchip (CUDA Toolkit 13.0.88, gcc 11.5.0) for $m=k=n$. Methods shown in red and italics achieve ZGEMM-level accuracy.
@@
Figure 11: Throughput performance (in TFLOPS) of double-precision complex matrix multiplication on NVIDIA B200 Blackwell GPU with Intel Xeon Platinum 8570 CPU (CUDA Toolkit 12.8.93, g++ 13.3.0) for $m=k=n$. Methods shown in red and italics achieve ZGEMM-level accuracy.
@@
Figure 12: Throughput performance (in TFLOPS) of double-precision complex matrix multiplication on NVIDIA GeForce RTX 5080 GPU with AMD Ryzen 7 3700X CPU (CUDA Toolkit 13.0.88, gcc 11.5.0) for $m=k=n$. Methods shown in red and italics achieve ZGEMM-level accuracy.
@@
Figure 13: Throughput performance (in TFLOPS) of double-precision complex matrix multiplication on AMD Instinct MI300X GPU with AMD EPYC 9534 CPU (ROCm 6.4.2, clang 19.0.0git) for $m=k=n$. Methods shown in red and italics achieve ZGEMM-level accuracy.
Figure 14: Speedups of single-precision complex matrix multiplication over native CGEMM for $m=k=n=16384$
Figure 15: Speedups of double-precision complex matrix multiplication over native ZGEMM for $m=k=n=16384$
@@
These performance characteristics correlate strongly with the architectural balance of arithmetic throughput on each GPU. The GH200 provides a ratio of $\text{FP64}:\text{FP32}:\text{INT8}=1:1:29.5$, indicating that INT8 units are more abundant than FP64 and FP32 units. The B200 exhibits an even more pronounced imbalance, $1:2:121.6$, where INT8 throughput overwhelmingly dominates the floating-point pipelines. This architectural skew directly amplifies the benefit of the proposed kernels, which replace large portions of floating-point arithmetic with INT8-intensive computation. In contrast, the MI300X offers a more moderate ratio of $1:1:16$, with INT8 throughput being moderately higher than floating-point throughput. As a result, the computational substitution enabled by our method yields only marginal acceleration on the MI300X, consistent with the limited speedups observed in the benchmarks. For the RTX 5080, the observed performance gains, especially the very large speedups in ZGEMM, suggest a much higher INT8 throughput relative to that of floating-point units. The architectural balance on the RTX 5080 corresponds to $\text{FP64}:\text{FP32}:\text{INT8}=1:64:512$, which explains the pronounced acceleration when using INT8-based emulation.
@@
Compared to the Ozaki-I emulation implemented in cuBLAS, the proposed methods delivered substantially higher throughput on both the GH200 and RTX 5080. On the GH200, Ozaki-I with 7–9 slices achieved 25–39 TFLOPS, whereas the proposed methods achieved 99–136 TFLOPS. On the RTX 5080, Ozaki-I achieved 6.4–10 TFLOPS, whereas the proposed methods achieved 19–26 TFLOPS. These disparities arise from fundamental algorithmic differences: Ozaki-I with $S$ slices requires $S(S+1)/2$ INT8 complex matrix multiplications, whereas the proposed technique with $N$ moduli requires only $N$ INT8 complex matrix multiplications, whose cost is further reduced by a factor of $3/4$ using the Karatsuba method. This reduction in both the number and cost of INT8 matrix multiplication kernel invocations explains the substantially higher performance of the proposed methods.
IV-C Supplemental Comparison of Real-Valued Emulation
@@
In this section, we focus on double-precision real matrix multiplication. We compare two emulation strategies: the Ozaki-I based emulation provided by cuBLAS and the Ozaki-II based emulation proposed in [uchino_ozaki2], enhanced with the blocking technique introduced in Section III. This supplemental analysis aims to isolate the behavior of Ozaki-type schemes in the real domain and to clarify the performance implications of incorporating blocking into Ozaki-II under conditions consistent with real-valued DGEMM emulation.
@@
According to [uchino_ozaki2], the fast mode and the accurate mode require 14–18 and 14–17 moduli, respectively, to achieve DGEMM-level accuracy. On the GH200 with matrix size $m=n=k=16384$, the Ozaki-II fast-$N$ configuration with $14\leq N\leq 18$ achieved 63–82 TFLOPS and the accu-$N$ configuration with $14\leq N\leq 17$ achieved 63–75 TFLOPS. Incorporating the blocking technique improved performance across all settings: the blocked fast-$N$ version achieved 72–93 TFLOPS and the blocked accu-$N$ version achieved 71–85 TFLOPS. cuBLAS’s Ozaki-I implementation with 7–10 slices achieved only 20–39 TFLOPS and the standard DGEMM kernel achieved 61 TFLOPS.
@@
On the RTX 5080 with $m=n=k=8192$, the fast-$N$ and accu-$N$ configurations of the proposed methods achieved 12–16 TFLOPS and 12–15 TFLOPS, respectively, and cuBLAS’s Ozaki-I achieved 5.4–10 TFLOPS. In comparison, the native DGEMM kernel achieved 0.84 TFLOPS, highlighting the substantial advantage of INT8-intensive Ozaki-type emulation on this consumer architecture.
@@
These results demonstrate that blocking substantially enhances the performance of Ozaki-II for large-scale problems, particularly on bandwidth-rich GPUs such as the B200, and that Ozaki-II, whether blocked or unblocked, consistently outperforms cuBLAS’s Ozaki-I due to its reduced number of INT8 matrix multiplication invocations.
@@ V Conclusion
This study proposed emulation methods for complex matrix multiplication based on the Ozaki-II scheme and developed a portable library that supports execution on both AMD and NVIDIA GPUs. The numerical results demonstrated that the proposed methods outperform the standard GEMM routines in cuBLAS and hipBLAS, which are implemented using native FP32 and FP64 arithmetic. In addition, the proposed methods outperform emulation using the Ozaki-I scheme implemented in cuBLAS.
Future work will include comparisons with the BF16x9 CGEMM emulation algorithm provided for NVIDIA B200 and B300 GPUs introduced in cuBLAS 12.9. A major limitation of the proposed methods, as well as emulation-based approaches in general, is the substantial working memory required. This overhead is currently unavoidable when aiming to achieve high performance without relying on FP32 or FP64 matrix multiplication. Addressing this issue remains an open challenge not only for emulation techniques but also for high-performance computing applications that employ them.
The Ozaki-I and Ozaki-II schemes utilize INT8 matrix engines, benefiting from the fact that the error-free low-precision matrix multiplications in these schemes do not require exponent handling, enabling efficient synergy with INT8 operations with INT32 accumulation. When using FP16 or BF16 matrix engines with FP32 accumulation, the effective representable integer range is limited to 24 bits due to the FP32 significand, leading to significantly reduced computational efficiency compared to that of INT32 accumulation. However, some recent GPUs, such as the NVIDIA B300, provide reduced INT8 capabilities. Support for such hardware architectures will be considered in future work.